### v1

@@ 5 Discussion
With regards to scalability beyond one million data points, [40] derived a formula to predict how much time it would take to calculate the covariance matrix for $N$ data points: $T=\frac{N^{2}t_{b}}{2nb^{2}}$. Here, $N$ is the size of the dataset, $t_{b}$ is the time it takes for a covariance batch to be computed, $n$ is the number of workers, and $b$ is the batch size. The validity of this formula has been well-tested in our experiments. Assuming $n=20,000$ available GPUs that can calculate a batch of size $b=15,000$ in 2 seconds, we could calculate the covariance matrix for 100 million data points in $2$ seconds; with the current algorithm, the log-determinant and the MINRES could be calculated in about 1000 seconds each. This results in a log marginal likelihood evaluation time of just over an hour. Whether or not this is still in the realm of feasibility depends on the total available computing time. If the number of hyperparameters is relatively small, the MCMC might converge after around 500 iterations, which amounts to just over 3 weeks of total computing time. This toy example reveals that a 100 million-data-point GP is on the horizon. Note, however, that this calculation assumes constant sparsity similar to the one observed in the real data application in Section 4.2. Early evidence suggests that larger datasets will allow the proposed kernel to discover more sparsity. While this will not affect the calculation of the covariance matrix, it vastly reduces linear system solutions and log-determinant calculations.
@@
    This research was supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Contract No. DE-AC02-05CH11231 at Lawrence Berkeley National Laboratory.
@@
Ethics approval and consent to participate. Not applicable. Consent to participate. Not applicable. Consent for publication. Not applicable.
Availability of data and materials. The in situ temperature records supporting this article are based on publicly available measurements from the National Centers for Environmental Information (ftp://ftp.ncdc.noaa.gov/pub/data/ghcn/daily/).
@@
Appendix A Proof of Proposition 1
@@
Scenario (A.2) and (A.3) both lead to terms $\partial^{k}g$ with $k>n$ or $\partial^{l-k}h$ with $l-k>m$. All terms exist and are continuous only for scenario (A.1), which concludes the proof. ∎
@@ Appendix B Parametric nonstationary kernel
$C_{\text{core}}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{\text{core}})=\sigma({\bf x})\sigma({\bf x}^{\prime})\frac{\big{(}\Sigma({\bf x})\Sigma({\bf x}^{\prime})\big{)}^{d/4}}{\left(\frac{\Sigma({\bf x}^{\prime})+\Sigma({\bf x}^{\prime})}{2}\right)^{d/2}}\mathcal{M}_{\nu}\left(\sqrt{Q({\bf x},{\bf x}^{\prime})}\right),\hskip 12.91663pt{\bf x},{\bf x}^{\prime}\in\mathcal{X},$ (B.1)
@@
$Q({\bf x},{\bf x}^{\prime})=\frac{({\bf x}-{\bf x}^{\prime})^{\top}({\bf x}-{\bf x}^{\prime})}{\frac{1}{2}[\Sigma({\bf x})+\Sigma({\bf x}^{\prime})]}=\frac{||{\bf x}-{\bf x}^{\prime}||^{2}}{\frac{1}{2}[\Sigma({\bf x})+\Sigma({\bf x}^{\prime})]},$
and $\mathcal{M}_{\nu}(\cdot)$ is the Matérn correlation function with smoothness $\nu$. In Eq. B.1, $\sigma(\cdot)$ is the signal standard deviation, and $\Sigma(\cdot)$ controls the length-scale of correlation. This kernel is locally isotropic; note that it is straightforward to generalize Eq. B.1 to be locally anisotropic [see, e.g., 45]. One can then parameterize $\sigma(\cdot)$ and $\Sigma(\cdot)$ to be log-linear functions of a set of basis functions, e.g.,
@@
$\phi_{(\cdot)}^{m}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,v_{(\cdot)})\hskip 12.91663pt\text{where}\hskip 12.91663ptv_{(\cdot)}\sim U(0,10^{5}).$
@@ Appendix C Gaussian process training via Markov chain Monte Carlo
(here, proportionality refers to quantities involving $\pi_{ij}$). Noting that the prior for $\pi_{ij}$ (uniform on the unit interval) can be re-written as a Beta(1,1)11(1,1)( 1 , 1 ) distribution, i.e.,
@@
which is proportional to a Beta$(1+a_{ij},2-a_{ij})$ distribution. Hence, the prior probabilities can be updated by sampling from a Beta$(1+a_{ij}^{\text{curr}},2-a_{ij}^{\text{curr}})$ distribution.
Appendix D Synthetic data examples
We now describe two synthetic data examples to demonstrate the performance of our proposed kernel under mis-specification of the number of bump functions (Section D.1) and relative to competing methods across a variety of ground-truth data states (Section D.2). In Section D.2, we assess performance quality using two metrics: (1) root mean square error (RMSE) for the test set, which assesses the pointwise error in posterior mean predictions; and (2) the continuous rank probability score [CRPS; 49] for the test set, which (unlike the RMSE) assesses both precision and sharpness of predictions. For example, the CRPS rewards a bad prediction with large uncertainty over a bad prediction with small uncertainty. Note that the reported CRPS scores are calculated for each test data point and then averaged across all test data.
D.1 Mis-specified bump functions
Figure D.1: Results for the synthetic data test described in Sections D.1. The ground truth function and noisy measurements are shown in panel (a); panels (b) and (c) show best estimates of the bump functions (posterior medians) and posterior predictions with uncertainty (posterior mean and 90% credible intervals), respectively. Results are shown when the kernel has too few bump functions (left), the correct number of bump functions (center), and too many bump functions (right).
and we have noisy measurements from this function for training a GP (see Figure D.1a.). The form of this ground truth function suggests that there should be two sets of two bump functions: $y(x)$ is perfectly correlated in $(0,0.25]$ and $(0.5,0.75]$; furthermore, $y(x)$ is perfectly correlated in $(0.25,0.5]$ and $(0.75,1)$. In other words, $n_{1}=2$ and $n_{2}=2$. We then fit our proposed kernel to the noisy data assuming we know nothing about its ground truth, under three configurations of bump functions: too few bump functions, where we set $n_{1}=1$ and $n_{2}=2$; the right number of bump functions, using $n_{1}=2$ and $n_{2}=2$; too many bump functions, using $n_{1}=4$ and $n_{2}=4$. Notice that although the second configuration is the only one with the correct number of bump functions, in all cases we infer the bump functions characteristics (positions and radii) directly from the data. When fitting, we suppose the core kernel is a stationary Matérn with smoothness $\nu=2.5$.
Results are shown in Figure D.1(b) and (c), where we show both the posterior median estimates of the bump functions and the posterior predictions of $y(\cdot)$ with uncertainty, respectively. First, consider the bump function estimates in Figure D.1(b). When there are too few bump functions, the training correctly identifies the position and size of the bump functions for part of the domain, i.e., that $(0.25,0.5]$ is correlated with $(0.75,1)$, while the kernel does not have the flexibility to capture the distance-unrelated correlations in $(0,0.25]$ and $(0.5,0.75]$. As such, the predictions capture the ground truth function with very low uncertainty in $(0.25,0.5]$ and $(0.75,1)$ but revert to the sparse stationary kernel in the other parts of the domain – effectively overfitting the data in $(0,0.25]$ and $(0.5,0.75]$. When there are the correct number of bump functions, the training correctly identifies where these bump functions should go and how large they should be; consequently, posterior predictions are correctly calibrated over the entire unit interval. Finally, when there are too many bump functions, the algorithm correctly determines both the size and position of the bump functions, but also that only four of the 16 total bump functions should be activated. The other 12 are appropriately zeroed out. Posterior predictions are effectively identical to those generated using the correct number of bump functions.
D.2 Comparison with competing methods
  1. S1
    Non-sparse and stationary: $C_{y}({\bf x},{\bf x}^{\prime})=\sigma^{2}\mathcal{M}_{\nu}(||{\bf x}-{\bf x}^{\prime}||)$, where $\mathcal{M}_{\nu}(\cdot)$ is the Matérn kernel with smoothness $\nu$.
  2. S2
    Sparse and stationary: $C_{y}({\bf x},{\bf x}^{\prime})=\sigma^{2}f_{0}({\bf x},{\bf x}^{\prime};r_{0})$, where $f_{0}$ is as in Eq. 2.1.
  3. S3
    Sparse and nonstationary: $C_{y}({\bf x},{\bf x}^{\prime})=\sigma({\bf x})\sigma({\bf x}^{\prime})f_{0}({\bf x},{\bf x}^{\prime};r_{0})$, where the signal variance $\sigma^{2}(\cdot)$ depends on the input locations.
  4. S4
    Non-sparse and nonstationary: here we take $C_{y}$ to be the nonstationary kernel described in [45].
Here we use S1-S4 to denote different “scenarios”; see Appendix E for more details on the data generating mechanisms and specific hyperparameters used. Sample draws from the true parent distribution for a one-dimensional input space are shown in Supplemental Figure E.1. In each case, the synthetic data drawn from the parent distribution represents a ground truth function; we can therefore generate $N_{\text{train}}=50$ (randomly sampled) and $N_{\text{test}}=300$ (regularly spaced) data points for validation. Each training data set will include the addition of white noise that is approximately 10% of the signal variance. For each scenario, we generate $N_{\text{rep}}=100$ replicates or draws from the ground truth distribution; note that the locations of the training data are different across each replicate.
For each scenario and each replicate, we fit a series of constant-mean GP models to the synthetic data, mirroring the example given in Figure 1 (the “M” denotes a statistical model):
  • M1
    Stationary Gaussian process with a Matérn kernel
  • M2
    Sparse general Vecchia (SGV) approximate GP method as proposed in [32]
  • M3
    The sparsity-discovering kernel proposed in Section 2, using a stationary Matérn for the core kernel
  • M4
    The sparsity-discovering kernel proposed in Section 2, using the nonstationary core kernel defined in Equation B.1 where the $s_{m}(\cdot)$ are natural smoothing splines [59] and $M_{\sigma}=M_{\Sigma}=4$.
We note that M1 and M2 could in principle involve nonstationary kernels; however, we maintain stationary kernels for two reasons. First, M1 is primarily used here as a reference for what one can obtain with a “default” (stationary and exact) GP when working with smaller data sets; the whole point of this methodology is that M1 will become unusable for much larger data sets. Second, all of the underlying theory for Vecchia approximations [32] has been derived for stationary kernels. It is, of course, possible to use the Vecchia framework with a nonstationary kernel [see, e.g., 45]; however, without further theoretical developments there are no guarantees that the subjective choices one must make for the Vecchia approximation (size of conditioning sets, ordering of points, selection of nearest neighbors) will result in an appropriate approximation for a nonstationary kernel. The different data-generating mechanisms and statistical models are summarized in Table D.1. Our implementation of M1 and M2 involve standard, off-the-shelf implementations of these methods as enabled by the BayesNSGP software package for R [60]. For M3 and M4, the main user-specified feature is determining how many bump functions are allowed: here, we set $n_{1}=4$ and $n_{2}=4$). In order to ease comparison, we show CRPS and RMSE for M2-M4 relative to corresponding quantities from M1: relative scores less than one indicate RMSE and CRPS are smaller (and therefore better) for the model of interest versus a stationary GP.
Table D.1: Summary of five data-generating mechanisms (or “scenarios”) used to generate synthetic data and four kernels used within a Gaussian process to fit each of $N_{\text{rep}}=100$ repeated draws from each scenario.
Data generating mechanisms Kernels within the Gaussian process
Label Description Label Description
S1 Non-sparse, stationary M1 Stationary Matérn kernel
S2 Sparse, stationary M2 Vecchia approximation with Matérn
S3 Sparse, nonstationary M3 Sparsity-discovering kernel with stationary core
S4 Non-sparse, nonstationary M4 Sparsity-discovering kernel with nonstationary core
Appendix E Synthetic data examples: hyperparameters and sample data sets

### v3

@@ 6 Discussion
With regards to scalability beyond one million data points, [50] derived a formula to predict how much time it would take to calculate the covariance matrix for $N$ data points: $T=\frac{N^{2}t_{b}}{2nb^{2}}$. Here, $N$ is the size of the dataset, $t_{b}$ is the time it takes for a covariance batch to be computed, $n$ is the number of workers, and $b$ is the batch size. The validity of this formula has been well-tested in our experiments. Assuming $n=20,000$ available GPUs that can calculate a batch of size $b=15,000$ in 2 seconds, we could calculate the covariance matrix for 100 million data points in 2222 seconds; with the current algorithm, the log-determinant and the MINRES could be calculated in about 1000 seconds each. This results in a log marginal likelihood evaluation time of just over an hour. Whether or not this is still in the realm of feasibility depends on the total available computing time. If the number of hyperparameters is relatively small, the MCMC might converge after around 500 iterations, which amounts to just over 3 weeks of total computing time. This toy example reveals that a 100 million-data-point GP is on the horizon. Note, however, that this calculation assumes constant sparsity similar to the one observed in the real data application in Section 5.2. Early evidence suggests that larger datasets will allow the proposed kernel to discover more sparsity. While this will not affect the calculation of the covariance matrix, it vastly reduces linear system solutions and log-determinant calculations.
@@
    This research was supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research’s Applied Mathematics program under Contract No. DE-AC02-05CH11231 at Lawrence Berkeley National Laboratory.
@@
Ethics approval and consent to participate. Not applicable.
Consent to participate. Not applicable.
Consent for publication. Not applicable.
Availability of data and materials. The in situ daily maximum temperature records supporting this article are based on publicly available measurements from the National Centers for Environmental Information (ftp://ftp.ncdc.noaa.gov/pub/data/ghcn/daily/).
@@
Appendix A Proof of Propositions
Proposition 1. The kernel $C_{\text{sparse}}$ is strictly positive definite. Furthermore, $C_{y}$ is strictly positive definite whenever $C_{\text{core}}$ is. More formally, for every admissible parameter vector
$\vartheta=\bigl(s_{0},r_{0},\{h_{ij},a_{ij},b_{ij},r_{ij}\}\bigr)\quad\text{with }s_{0}>0,$
the kernel
$C_{\text{sparse}}(x,x^{\prime};\vartheta)=s_{0}\,f_{0}\!\bigl(x,x^{\prime};r_{0}\bigr)+\sum_{i=1}^{n_{1}}\Bigl[\sum_{j=1}^{n_{2}}g_{ij}(x;\eta_{ij})\Bigr]\,\Bigl[\sum_{j=1}^{n_{2}}g_{ij}(x^{\prime};\eta_{ij})\Bigr]$ (A.1)
is strictly positive-definite (PD) on every finite subset of $\mathcal{X}\subset\mathbb{R}^{d}$.
Proof.
For each fixed $i$ define
$\phi_{i}(x)=\sum_{j=1}^{n_{2}}g_{ij}\bigl(x;\eta_{ij}\bigr),\qquad K_{i}(x,x^{\prime})=\phi_{i}(x)\,\phi_{i}(x^{\prime}).$
Given points $x_{1},\dots,x_{m}$, the Gram matrix $K_{i}(x_{p},x_{q})]_{pq}$ equals $\Phi_{i}\Phi_{i}^{\top}$ with $\Phi_{i}=(\phi_{i}(x_{1}),\dots,\phi_{i}(x_{m}))^{\top}$ and is therefore PSD; see Section 2.2 of [22]. Because the sum of PSD kernels is PSD [Proposition 2.3.3 of 62],
$K_{\text{bump}}(x,x^{\prime})=\sum_{i=1}^{n_{1}}K_{i}(x,x^{\prime})$
is PSD for all choices of $\eta_{ij}=(h_{ij},a_{ij},b_{ij},r_{ij})$. Then we come to examine the compactly supported function $f_{0}(x,x^{\prime};r_{0})=\varphi\!\bigl(\|x-x^{\prime}\|/r_{0}\bigr)$ with $\varphi$ as in [Theorem 4.9 36], which is strictly PD on $\mathbb{R}^{d}$. It is clear that multiplying by $s_{0}>0$ preserves strictness.
If $K_{1}$ is strictly PD and $K_{2}$ is PSD, then $K_{1}+K_{2}$ is strictly PD because $v^{\top}(K_{1}+K_{2})v\geq v^{\top}K_{1}v>0$ for any non-zero vector $v$. Applying this to $s_{0}f_{0}$ and $K_{\text{bump}}$ shows that $C_{\text{sparse}}$ in (A.1) is strictly PD. So far we require only $s_{0},r_{0},r_{ij}>0$; no further constraints are placed on $h_{ij},a_{ij},b_{ij}$. Consequently, every parameter draw encountered during optimization or MCMC, and every prediction location, retains strict positive-definiteness. The model uses the Hadamard product $C_{y}(x,x^{\prime})=C_{\text{core}}(x,x^{\prime};\theta_{\text{core}})\,C_{\text{sparse}}(x,x^{\prime};\vartheta)$. By the Schur product theorem [Theorem 5.2.1 of 41] the element-wise product of positive- definite matrices is positive-definite, so $C_{y}$ is strictly PD whenever $C_{\text{core}}$ is. ∎
@@ Appendix A Proof of Propositions
Scenario (A.3) and (A.4) both lead to terms $\partial^{k}g$ with $k>n$ or $\partial^{l-k}h$ with lk>ml-k>m. All terms exist and are continuous only for scenario (A.2), which concludes the proof. ∎
A.1 Fixed-domain asymptotics for the bump-product kernel
Notation follows the manuscript exactly: $\mathcal{X}\subset\mathbb{R}^{d}$ is bounded and convex; the infill design places points $\mathcal{D}_{N}=\{x_{1},\dots,x_{N}\}\subset\mathcal{X}$ with minimum spacing $\delta_{N}=\min_{i\neq j}\|x_{i}-x_{j}\|\to 0$ as $N\to\infty$. The data-generating Gaussian process is
$y(x)\sim\mathrm{GP}\!\bigl(0,\;C_{y}^{\star}(x,x^{\prime})\bigr),\qquad C_{y}^{\star}=C_{\text{core}}(x,x^{\prime};\theta_{\text{core}}^{\star})\,C_{\text{sparse}}(x,x^{\prime};\theta_{\text{sparse}}^{\star}),$
where $C_{\text{sparse}}$ has the bump-product form of Eq. (3) in the paper. All hyper-parameters $\theta_{y}=(\theta_{\text{core}},\theta_{\text{sparse}})$ receive priors with positive, continuous density in a neighbourhood of $\theta_{y}^{\star}$.
Proposition 3.
For every $\varepsilon>0$, assume that the prior of $\theta_{y}$ have probability mass to every neighborhood of $\theta_{y}^{\star}$ then its posterior
$\Pi_{N}\bigl(\|\theta_{y}-\theta_{y}^{\star}\|>\varepsilon\bigr)\;\xrightarrow[N\to\infty]{P_{\theta_{y}^{\star}}}\;0,$
where $\Pi_{N}$ is the posterior of $\theta_{y}$ given $\bigl(y(x_{1}),\dots,y(x_{N})\bigr)$.
Theorem 2.1 of [48] gives posterior contraction whenever (i) the likelihood is continuous in the finite-dimensional parameter, (ii) the prior assigns positive mass to every neighbourhood of the truth, and (iii) the model is identifiable locally. Each covariance entry is a smooth function of $\theta_{y}$, so the likelihood is continuous; the prior requirement is built in; Proposition 2 of our current paper guarantees local identifiability. Although the bump radii $r_{ij}$ constitute a varying support, they lie in a fixed compact set and the prior places mass near $r_{ij}^{\star}$. Hence Theorem 2.1 applies verbatim.
Proposition 4 (stabilisation of the sparsity pattern).
Let
$r_{\max}^{\star}=\max_{0\leq i\leq n_{1},\,1\leq j\leq n_{2}}r_{ij}^{\star},\qquad\delta_{N}=\min_{i\neq j}\|x_{i}-x_{j}\|,\qquad\Omega_{N}(\theta_{y})=\bigl[C_{y}(x_{i},x_{j};\theta_{y})\bigr]_{i,j=1}^{N}.$
For any deterministic sequence $\varepsilon_{N}\downarrow 0$ define
$K_{N}(\varepsilon_{N})\;=\;\frac{\text{Vol}\!\bigl(B_{d}(2(r_{\max}^{\star}+\varepsilon_{N})+\tfrac{\delta_{N}}{2})\bigr)}{\text{Vol}\mathcal{}\!\bigl(B_{d}(\tfrac{\delta_{N}}{2})\bigr)}\;=\;\Bigl(\tfrac{4(r_{\max}^{\star}+\varepsilon_{N})}{\delta_{N}}+1\Bigr)^{d}.$
Let
$\mathcal{A}_{N}=\Bigl\{\theta_{y}:\,\max_{i,j}|r_{ij}-r_{ij}^{\star}|\leq\varepsilon_{N}\Bigr\}.$
Then
[…]
where $\Omega_{N}(\theta_{y})=\bigl[C_{y}(x_{i},x_{j};\theta_{y})\bigr]_{i,j=1}^{N}$.
Proof. Proposition 3 implies $\max_{i,j}|r_{ij}-r_{ij}^{\star}|=o_{P}(1)$. Hence, with high posterior probability, the support of $C_{\text{sparse}}$ is contained in balls of radius $r_{\max}^{\star}+M_{N}\delta_{N}$. A volume-packing argument bounds the number of design points per ball, giving the stated uniform row sparsity.
Theorem 1 gives posterior consistency of every bump radius, so $\Pi_{N}(\mathcal{A}_{N})\to 1$. We fix $\theta_{y}\in\mathcal{A}_{N}$ and set $R_{N}=2(r_{\max}^{\star}+\varepsilon_{N})$. If $C_{\text{sparse}}(x_{i},x_{j};\theta_{y})\neq 0$, then either
$f_{0}(x_{i},x_{j};r_{0})\neq 0\quad\text{or}\quad g_{ij}(x_{i};\cdot)\,g_{ij}(x_{j};\cdot)\neq 0.$
In both cases $\|x_{i}-x_{j}\|\leq R_{N}$, because $r_{0}$ and every $r_{ij}$ lie in $0,r_{\max}^{\star}+\varepsilon_{N}]$. Next we can give a packing argument, and we fix $i$ and set
$\mathcal{N}_{i}(\theta_{y})=\bigl\{j:\,C_{\text{sparse}}(x_{i},x_{j};\theta_{y})\neq 0\bigr\}\subset B_{d}(x_{i},R_{N}).$
Because the design is $\delta_{N}$-separated, the closed balls $B_{d}(x_{j},\delta_{N}/2)$, $j\in\mathcal{N}_{i}(\theta_{y})$, are pairwise disjoint and all lie inside $B_{d}(x_{i},R_{N}+\delta_{N}/2)$. Hence
$\mathcal{N}_{i}(\theta_{y})\;\text{Vol}\bigl(B_{d}(\tfrac{\delta_{N}}{2})\bigr)\;\leq\;\text{Vol}\bigl(B_{d}(R_{N}+\tfrac{\delta_{N}}{2})\bigr)\quad\Longrightarrow\quad\#\mathcal{N}_{i}(\theta_{y})\leq K_{N}(\varepsilon_{N}).$
On $\mathcal{A}_{N}$ every row of $\Omega_{N}(\theta_{y})$ has at most $K_{N}(\varepsilon_{N})$ non-zeros; since $\Pi_{N}(\mathcal{A}_{N})\to 1$, the corollary follows.
The classical fixed-support proofs (van der Vaart & van Zanten, Furrer et al., Stein) require only that the taper radius be finite and non-shrinking. Proposition 4 shows that the data-driven bump radii concentrate on $\{r_{ij}^{\star}\}$, so for large $N$ the support radius is stochastically constant. Conditioning on this high-probability event allows the existing theorems to be invoked unchanged; the proposed kernel therefore inherits the full suite of classical fixed-domain guarantees while discovering its own sparsity pattern from the data.
@@ Appendix B Parametric nonstationary kernel
$C_{\text{core}}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{\text{core}})=\sigma({\bf x})\sigma({\bf x}^{\prime})\frac{\big(\Sigma({\bf x})\Sigma({\bf x}^{\prime})\big)^{d/4}}{\left(\frac{\Sigma({\bf x}^{\prime})+\Sigma({\bf x}^{\prime})}{2}\right)^{d/2}}\mathcal{M}_{\nu}\left(\sqrt{Q({\bf x},{\bf x}^{\prime})}\right),\qquad{\bf x},{\bf x}^{\prime}\in\mathcal{X},$ (B.1)
@@
[…]
and $\mathcal{M}_{\nu}(\cdot)$ is the Matérn correlation function with smoothness $\nu$. In Eq. B.1, $\sigma(\cdot)$ is the signal standard deviation, and $\Sigma(\cdot)$ controls the length-scale of correlation. This kernel is locally isotropic; note that it is straightforward to generalize Eq. B.1 to be locally anisotropic [see, e.g., 46]. One can then parameterize $\sigma(\cdot)$ and $\Sigma(\cdot)$ to be log-linear functions of a set of basis functions, e.g.,
@@
$\phi_{(\cdot)}^{m}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,v_{(\cdot)})\qquad\text{where}\qquad v_{(\cdot)}\sim U(0,10^{5}).$
@@ Appendix C Gaussian process training via Markov chain Monte Carlo
(here, proportionality refers to quantities involving $\pi_{ij}$). Noting that the prior for $\pi_{ij}$ (uniform on the unit interval) can be re-written as a Beta(1,1)(1,1) distribution, i.e.,
@@
[…]
Appendix D Synthetic data examples: hyperparameters and sample data sets